# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ 3 Methods: Weighted Rolling Validation
$RV_{i}^{(k)}=RV\left(\left\{\hat{f}_{l}^{(k)},l\in[i]\right\},\xi\right)=\sum_{l=1}^{i}l^{\xi}\left(\hat{f}_{l}^{(k)}\left(X_{l+1}\right)-Y_{l+1}\right)^{2},$ (13)
where $\xi\geq 0$ is a rolling weight exponent to be chosen by the user. We will discuss some rule-of-thumb choices for it soon in Section 3.1. A positive $\xi$ will put more weight on larger values of $l$, which can improve finite-sample performance by alleviating some “detection delay” due to the cumulative nature of wRV. Although the definition of wRV statistics used squared-loss, the method is general and can be extended to other loss functions (see Section 8 for a quantile-regression example).
At each step $i$, we can use $\hat{k}^{*}_{i}=\arg\min_{k\in[K]}RV_{i}^{(k)}$ as an estimate of $k^{*}_{i}$ (9) or $k^{*}$ (11). A statistically favorable procedure is expected to achieve $P(\hat{k}^{*}_{i}=k^{*})=1-o(1)$ when the superior sequence exists.
It is direct to see that $RV_{i}^{(k)}$ is an unbiased estimator of accumulated prediction error:
$E\left[RV_{i}^{(k)}\right]=\sum_{l=1}^{i}l^{\xi}E\left[\left(\hat{f}_{l}^{(k)}(X)-Y\right)^{2}\right].$ (14)
@@
    There is a gap between $E\left[RV_{i}^{(k^{*})}\right]$ and $\inf_{k\neq k^{*}}E\left[RV_{i}^{(k)}\right]$, which is closely related to the excess risk condition (11).
@@
    The sample versions $RV_{i}^{(k)}$ do not deviate too far away from their expected values (specifically it should be smaller than the aforementioned “gap”). This is technically more challenging as the dependence structure in $RV_{i}^{(k)}$ is more complicated than simple martingales. We resort to a stability-based argument to formally quantify its variability.
@@ 3.1 Choice of Weighting Exponent
A positive weighting exponent $\xi$ can significantly enhance finite-sample model selection quality, which is also demonstrated later in the numerical examples (Section 7). The optimal $\xi$ depends on the specific problem and the estimators utilized, making it challenging to determine in practice. In general, we recommend using a fixed $\xi=1$ to balance selection sensitivity and stability. This choice is determined under certain assumptions on the excess risks. We provide some intuition below and the technical details can be found in Appendix F.
@@
Assume we receive a sequence of IID samples $(X_{i},Y_{i})=(X_{i},f_{0}(X_{i})+\epsilon_{i})$, where $\epsilon_{i}$ is a centered noised variable $E[\epsilon_{i}\mid X_{i}]=0$. Consider a sequences of estimators $\{f_{i}^{(1)}\}$ whose risk at step $i$ is
$E\left[\left(f_{i}^{(1)}(X)-Y\right)^{2}\right]=E[r_{i,1}]+E\left[\epsilon_{i}^{2}\right].$
@@
We similarly denote the excess risk of another sequence $\{f_{i}^{(2)}\}$ as $E[r_{i,2}]$. After revealing $i$ samples, ideally, we want to instantly identify the one with a smaller risk and use it for predictions or downstream procedures. However, wRV only provides a noisy version of accumulative risks $E\left[RV_{i}^{(k)}\right]$, $k=1,2$, whose relative orders may differ from that of $E[r_{i,k}]$ for finite $i$.
@@
To simplify the discussion, assume both models have polynomial orders of risk: $E[r_{i,1}]=Ai^{-a}$ and $E[r_{i,2}]=Bi^{-b}$ with $0<A<B$ and $10\leq a<b<$ 1. For smaller sample sizes $i\leq(B/A)^{1/(b-a)}$, $\{f_{i}^{(1)}\}$ is a better estimator, but their relative performance flips afterward because $f_{i}^{(2)}$ converges to $f_{0}$ at a faster rate. However, the relative magnitudes of $E\left[RV_{i}^{(1)}\right]$ and $E\left[RV_{i}^{(2)}\right]$ do not switch until $i\geq(B/A)^{1/(b-a)}T(a,b,\xi)$ with a delay ratio $T(a,b,\xi)>1$ (whose value interestingly does not depend on $A$ or $B$).
@@
A calculation (Lemma F.2) shows that a larger weighting exponent $\xi$ can reduce the delay ratio $T(a,b,\xi)$. Specifically, given a $t>1$, we can roughly control $T(a,b,\xi)$ to be smaller than $t$ when using a $\xi>(\log t)^{-1}$, illustrated in Figure F.1. To enhance the algorithm’s sensitivity, a larger $\xi$ that minimizes the delay ratio is preferred. However, this comes at the cost of reducing the concentration of $RV_{i}^{(k)}$ on $E[RV_{i}^{(k)}]$ by a constant factor. For example, if $\xi=3$, the $i$-th sample has a weight of $i^{3}$ when calculating $RV_{i}$, while the $i/2$-th sample has a weight of $i^{3}/8$, contributing much less to the wRV criterion. We recommend selecting $\xi=1$ to control the worst-case $T(a,b,\xi)$ at no more than than $2.4$ while maintaining good stability.
@@
We can also compare the excess risk of the selected sequence to that of the optimal one. When $i$ is slightly smaller than $I_{F}=(B/A)^{1/(b-a)}T(a,b,\xi)$, the difference between $E[r_{1,i}]$ and $E[r_{2,i}]$ is maximized, while wRV on average still prefer the sub-optimal sequence, as discussed above. We can compute $E[r_{1,I_{F}}]/E[r_{2,I_{F}}]$ and use it as a measure of the worst-case efficiency loss. It takes a simple form
$E[r_{1,I_{F}}]/E[r_{2,I_{F}}]=1+\frac{b-a}{\xi+1-b}\leq 1+\xi^{-1}.$ (15)
Choosing $\xi=1$ controls the efficiency loss within twice that of the optimal choice.
@@
The wRV-based decision improves dramatically when $\xi$ takes a positive value but stabilizes after it is larger than $1$. We illustrate this phenomenon in Section 7 and Section H.2—the ranking plots for $\xi=0$ differ significantly from the rest.
@@ 3.2 Computational Advantages of Rolling Validation
Input: A stream of training samples $(X_{i},Y_{i}),i\in\mathbb{N}^{+}$. The estimator updating rule $\text{Update}(f,(X,Y),\lambda)$. Candidate hyperparameter sequences $\{\lambda^{(k)}_{i}:i\in\mathbb{N}^{+}\}$, $k\in[K]$, weight exponent $\xi$.
Initialize Estimators $\hat{f}^{(1)}=...=\hat{f}^{(K)}=0$. wRV statistics $RV^{(1)}=...=RV^{(K)}=0$. Training sample index $i=0$.
while there are training samples not processed do
       Read in a new training sample $(X_{i+1},Y_{i+1})$
for $k$ in $1$ to $K$ do
             Calculate prediction from the $k$th estimate $\hat{Y}_{i+1}\leftarrow\hat{f}^{(k)}(X_{i+1})$
Update its wRV statistics $RV^{(k)}\leftarrow RV^{(k)}+i^{\xi}(Y_{i+1}-\hat{Y}_{i+1})^{2}$
Update the estimator $\hat{f}^{(k)}\leftarrow\text{Update}(\hat{f}^{(k)},(X_{i+1},Y_{i+1}),\lambda^{(k)}_{i+1})$
@@
      Output current selection $\hat{k}_{i}^{*}=\arg\min_{k}RV^{(k)}$
Increase training sample index $i\leftarrow i+1$
@@
Algorithm 1 Stochastic approximation with wRV for regression problems.
The wRV statistics (13) is not only a natural target to examine when comparing online estimators, but it can also save much computational expense when implemented efficiently. In Algorithm 1, we summarize our recommended framework of integrating model training and model selection in an online fashion. In contrast to the batch-learning scenario where training, validation, and testing are typically done in separated phases, we proposed a procedure that: 1) updates multiple estimators simultaneously; 2) calculates the wRV statistics interchangeably with model training; and 3) can offer a selected model at any time during the training process.
We take the kernel-SGD estimator (5) as an example to explicitly quantify the computational gain compared with direct CV. Examining the updating rule, we can verify that both the trajectory functions $\{\tilde{f}_{i}\}$ and the estimator $\{\hat{f}_{i}\}$ are a linear combination of $i$ kernel functions “centered” at the past sample covariate vectors. Formally, there are some $\tilde{\beta}_{mi},\hat{\beta}_{mi}\in\mathbb{R}:$
@@
Suppose evaluating one kernel function $\mathcal{K}(x,z)$ takes $O(p)$ computation—recall that $p$ is the dimension of $X_{i}$. Given a new sample $(X_{i+1},Y_{i+1})$, the majority of computation for the Update operator is calculating $\mathcal{K}(X_{i+1},X_{m})$ for all $m\in[i]$. This means $O((p+C)i)$ calculation when processing one sample $(X_{i+1},Y_{i+1})$, where $C$ is a small constant corresponding to the extra computation after evaluating all $\mathcal{K}(X_{i+1},X_{m})$. Moreover, the above computational expense accumulates to the order of $O((p+C)n^{2})$ for processing the first $n$ samples $\{(X_{1},Y_{1}),...,(X_{n},Y_{n})\}$. Naively training $K$ kernel-SGD estimators one at a time with $n$ samples would require $O(K(p+C)n^{2})$ computation if no intermediate quantities are saved. However, when the $K$ sequences of trajectories only differ at the learning rate $\gamma_{i}$, they can be simultaneously updated with only $O((p+KC)n^{2})$ computation using the scheme in Algorithm 1, simply because the evaluated $\mathcal{K}(X_{i+1},X_{m})$ kernel value at each step can be shared between the estimators.
In addition to the computational gain for simultaneous training, these evaluated kernel values can be further shared with the wRV statistic calculation that only increases the total expense to at most $O((p+2KC)n^{2})$. This basically means that performance assessment is a “free lunch” compared with the expense of training, without the need for a hold-out data for model selection. The only extra space expense is saving extra $K$ numbers $\{RV^{(k)}\}$. A similar calculation also holds for Sieve-SGD (6): the computational expense of training and validating $K$ estimators can be bounded by $O((p+2KC)nJ_{n})$, where $J_{n}$ is the largest basis number among the $K$ candidate estimators. For a fixed $p$, $J_{n}$ is typically a diverging sequence with $J_{n}=o(\sqrt{n})$.
@@ 4 Consistency of Rolling Validation, Fixed Number of Candidates
In this section, we investigate the statistical properties of the proposed RV procedure for regression problems under the squared loss. We focus on whether wRV can select $k^{*}$ specified in (11) from a finite candidate set—the increasing candidate-cardinality case is covered in Section 5. We will show that when the online generalization errors of these estimators are sufficiently different and the wRV risk estimates are stable under small perturbations of the training data, wRV can select the better model with probability approaching one. For simplicity, we consider $K=2$ with two sequences of online estimators of the conditional mean function $f_{0}$. This consistency result extends to any finite $K\geq 2$ by considering $k^{*}$ against each candidate and using the union bound. Below, we list the technical conditions.
@@
The data points $\{(X_{i},Y_{i})\in\mathbb{R}^{p+1},i\in\mathbb{N}^{+}\}$ are IID samples from a common distribution $P(X,Y)$. The centered noise variable $\epsilon_{1}=Y_{1}-f_{0}(X_{1})$ has finite variance: $E[\epsilon_{1}^{2}|X_{1}=x]\leq\sigma^{2}$ for any $x$ in the support of $X_{1}$.
@@
Next, we formalize the assumption on the sub-optimal model’s estimation quality, essentially stating it is no better than $Mi^{-a}$ on average.
@@
$\liminf_{i\rightarrow\infty}i^{a}E[r(\hat{f}_{i})]=M.$ (16)
Recall $r(\hat{f}_{i})=$ $E_{X_{0}}\left[\left\{\hat{f}_{i}(X_{0})-f_{0}(X_{0})\right\}^{2}\right]$ and the outer expectation is unconditional, making the left-hand-side in (16) non-random. We impose a similar condition for the superior sequence, flipping the limit inferior to limit superior:
@@
$\limsup_{i\rightarrow\infty}i^{a}E[r(\hat{f}_{i})]=M.$ (17)
@@
$\sup_{i}E[r(\hat{f}_{i})]<\infty$ and there exist constants $a\in[0,1]$, $0<C_{1}<\infty$ such that
$\limsup_{i\rightarrow\infty}i^{2a}E\left[\left\{\hat{f}_{i}\left(X\right)-f_{0}\left(X\right)\right\}^{4}\right]=C_{1}.$ (18)
@@
Although 4.4 implies 4.3 with $M=C_{1}^{1/2}$, we state these conditions separately because, in Theorem 4.1, the superior and inferior sequences are assumed to satisfy different combinations of them. We will compare the constants in (16) and (17), while the constant $C_{1}$ in (18) is of lesser importance. Additionally, we do not require the sequence $i^{a}E[r(\hat{f}_{i})]$ itself to have a limit as $i\rightarrow\infty$.
@@
The values of $(M,a)$ reflect the quality of the estimator, which partially reflects the problem’s difficulty. They are determined by the choice of fitting methods and/or tuning parameters. For optimal parametric SGD, $a=1$. For a minimax-optimal estimation procedure in a Sobolev ellipsoid (7)—as well as some related RKHS space—$a=2s/(2s+1)$. The constant $M$ is less explicit but usually depends on the axis lengths of the ellipsoid, interpreted as the Sobolev/RKHS norm of $f_{0}$ or as a measure of $f_{0}$’s smoothness. Mean-squared error (MSE) convergence results for online (nonparametric) estimators are available in the literature, including [10, 42, 39, 40, 9]. Specifically, [42] also provides an interesting lower bound for the MSE error.
In addition to the moment conditions above, we also need the following stability conditions to control certain “non-standard” deviation quantities that cannot be fully reduced to martingale-type concentration. Let $(X^{\prime}_{j},Y^{\prime}_{j})$ be an IID copy of $(X_{j},Y_{j})$ for a $j\geq 1$. The actual estimator $\hat{f}_{i}(x)$ is trained with samples $Z_{i}=\{(X_{l},Y_{l}),l\in[i]\}$ and can be explicitly expressed as $\hat{f}_{i}(x)=\hat{f}_{i}(x;Z_{i})$. We denote an imaginary dataset $Z_{i}^{j}$, obtained by replacing $(X_{j},Y_{j})$ in $Z_{i}$ by its IID copy $(X^{\prime}_{j},Y^{\prime}_{j})$. The imaginary estimator trained with $Z_{i}^{j}$ is $\hat{f}_{i}(x;Z_{i}^{j})$. We need conditions on the difference between $\hat{f}_{i}(x;Z_{i})$ and its perturb-one version: $\nabla_{j}\hat{f}_{i}\left(x\right):=\hat{f}_{i}\left(x;Z_{i}\right)-\hat{f}_{i}\left(x;Z_{i}^{j}\right)$.
@@
$i^{2b}E[\{\nabla_{j}\hat{f}_{i}(X)\}^{2}|F^{j}]\leq C_{2}\qquad\text{almost surely},$ (19)
where $F^{j}$ is the $\sigma$-algebra generated by the first $j$ samples in $Z_{i}$.
@@
The sup-norm stability condition in (19) can be simplified and relaxed to $E[\{\nabla_{j}\hat{f}_{i}(X)\}^{2}]\lesssim i^{-2b}$ if we strengthen the convergence condition in 4.4 to the sup norm: $\sup_{x}\{|\hat{f}_{i}(x)-f_{0}(x)|\}\lesssim i^{-a}$. In fact, we require the product $E[\nabla_{j}\hat{f}_{i}(X)\{\hat{f}_{i}(X)-f_{0}(X)\}\mid F^{i}]$ to be small enough in the proof (equation (55)), which allows for Hölder’s-inequality type trade-off between the stability term and the estimation error.
@@
$\lim_{n\rightarrow\infty}\operatorname{pr}\left[{\rm RV}\left(\left\{g_{i},i\in[n]\right\},\xi\right)<{\rm RV}\left(\left\{h_{i},i\in[n]\right\},\xi\right)\right]=1$

### v4

@@ 3 Methods: Weighted Rolling Validation
${\rm RV}_{i}^{(k)}={\rm RV}\left(\left\{\hat{f}_{l}^{(k)},l\in[i]\right\},\xi\right)=\sum_{l=0}^{i}l^{\xi}\left(\hat{f}_{l}^{(k)}\left(X_{l+1}\right)-Y_{l+1}\right)^{2},$ (9)
where $\xi\geq 0$ is a weighting exponent to be chosen by the user. We will discuss some rule-of-thumb choices for it soon in Section 3.1. A positive $\xi$ will put more weight on larger values of $l$—i.e., the samples processed later—which can improve finite-sample performance by alleviating some “detection delay” due to the cumulative nature of wRV. Although the definition of wRV statistics used squared loss, the method is general and can be extended to other loss functions (see Section 8 for a quantile-regression example).
At each step $i$, we can use $\hat{k}^{*}_{i}=\arg\min_{k\in[K]}{\rm RV}_{i}^{(k)}$ as an estimate of $k^{*}_{i}$ (7) or $k^{*}$ (8). A statistically favorable procedure is expected to achieve $\mathbb{P}(\hat{k}^{*}_{i}=k^{*})=1-o(1)$ when the superior sequence exists.
It is direct to see that ${\rm RV}_{i}^{(k)}$ is an unbiased estimator of accumulated prediction error:
$\mathbb{E}\left[{\rm RV}_{i}^{(k)}\right]=\sum_{l=1}^{i}l^{\xi}\mathbb{E}\left[\left(\hat{f}_{l}^{(k)}(X)-Y\right)^{2}\right].$ (10)
@@
    There is a gap between $\mathbb{E}\left[{\rm RV}_{i}^{(k^{*})}\right]$ and $\inf_{k\neq k^{*}}\mathbb{E}\left[{\rm RV}_{i}^{(k)}\right]$, which is closely related to the excess risk condition (8).
@@
    The sample quantities ${\rm RV}_{i}^{(k)}$ do not deviate too far away from their expected values—specifically, their deviation should remain smaller than the aforementioned ”gap.” This is technically more challenging as the dependence structure in ${\rm RV}_{i}^{(k)}$ is more complicated than simple martingales. We resort to a stability-based argument to formally quantify its variability.
@@ 3.1 Choice of Weighting Exponent
A positive weighting exponent $\xi$ can significantly enhance finite-sample model selection quality, which is also demonstrated later in the numerical examples (Section 7). The optimal $\xi$ depends on the specific problem and the estimators utilized, making it challenging to determine in practice. In general, we recommend using a fixed $\xi=1$ to balance selection sensitivity and stability. This choice is determined under certain assumptions on the excess risks. We provide some intuition below, and the technical details can be found in Appendix F.
@@
Assume we receive a sequence of IID samples $(X_{i},Y_{i})=(X_{i},f_{0}(X_{i})+\epsilon_{i})$, where $\epsilon_{i}$ is a centered noised variable $\mathbb{E}[\epsilon_{i}\mid X_{i}]=0$. Consider a sequences of estimators $\{f_{i}^{(1)}\}$ whose risk at step $i$ is
$\mathbb{E}\left[\left(f_{i}^{(1)}(X)-Y\right)^{2}\right]=\mathbb{E}[r_{i,1}]+\mathbb{E}\left[\epsilon_{i}^{2}\right].$
@@
We similarly denote the excess risk of another sequence $\{f_{i}^{(2)}\}$ as $\mathbb{E}[r_{i,2}]$. After revealing $i$ samples, ideally, we want to instantly identify $k_{i}^{*}$ and use it for predictions or downstream procedures. However, wRV only provides a noisy version of cumulative risks $\mathbb{E}\left[{\rm RV}_{i}^{(k)}\right]$, $k=1,2$, whose relative orders may differ from that of $\mathbb{E}[r_{i,k}]$ for finite $i$.
@@
To simplify the discussion, we assume both models have polynomial orders of excess risks: $\mathbb{E}[r_{i,1}]=Ai^{-a}$ and $\mathbb{E}[r_{i,2}]=Bi^{-b}$ with $0<A<B$ and $10\leq a<b<$ 1. For smaller sample sizes $i\leq(B/A)^{1/(b-a)}$, $\{f_{i}^{(1)}\}$ is a better estimator, but their relative performance flips afterward because $f_{i}^{(2)}$ converges to $f_{0}$ at a faster rate. However, the relative magnitudes of $\mathbb{E}\left[{\rm RV}_{i}^{(1)}\right]$ and $\mathbb{E}\left[{\rm RV}_{i}^{(2)}\right]$ do not switch until $i\geq(B/A)^{1/(b-a)}T(a,b,\xi)$ with a delay ratio $T(a,b,\xi)>1$ (whose value interestingly does not depend on $A$ or $B$).
@@
A calculation (Lemma F.2) shows that increasing the weighting exponent $\xi$ can reduce the delay ratio. Specifically, for any target threshold $t>1$, the maximum delay ratio $T(a,b,\xi)$ across all $(a,b)$ can be controlled to stay below $t$ by choosing $\xi>1/\log t$, as illustrated in Figure F.1. To enhance the algorithm’s sensitivity, a larger $\xi$ that minimizes the delay ratio is preferred. However, this comes at the cost of reducing the concentration of ${\rm RV}_{i}^{(k)}$ on $\mathbb{E}[{\rm RV}_{i}^{(k)}]$ by a constant factor. For example, if $\xi=3$, the $i$-th sample has a weight of $i^{3}$ when calculating ${\rm RV}_{i}$, while the $i/2$-th sample has a weight of $i^{3}/8$, contributing much less to the wRV criterion. We recommend selecting $\xi=1$ to control the worst-case $T(a,b,\xi)$ at no more than $2.4$ while maintaining good stability.
@@
We can also compare the excess risk of the selected sequence to that of the optimal one. When $i$ is slightly smaller than $I_{F}=(B/A)^{1/(b-a)}T(a,b,\xi)$, the difference between $\mathbb{E}[r_{i,1}]$ and $\mathbb{E}[r_{i,2}]$ is maximized, while wRV on average still prefer the sub-optimal sequence, as discussed above. We can compute $\mathbb{E}[r_{I_{F},1}]/\mathbb{E}[r_{I_{F},2}]$ and use it as a measure of the worst-case efficiency loss. It takes a simple form
$\mathbb{E}[r_{I_{F},1}]/\mathbb{E}[r_{I_{F},2}]=1+\frac{b-a}{\xi+1-b}\leq 1+\xi^{-1}.$
A choice of $\xi=1$ limits the efficiency loss to within a factor of two of the optimal.
@@
The wRV-based decision improves dramatically when $\xi$ takes a positive value, but the benefit stabilizes after it is larger than $1$. We illustrate this phenomenon in Section 7 and Section H.2—the ranking plots for $\xi=0$ differ significantly from the rest.
@@ 3.2 Computational Advantages of Rolling Validation
Input :
A stream of training samples $(X_{i},Y_{i})$ for $i\in\mathbb{N}^{+}$.
An estimator update rule: $\text{Update}(f,(X,Y),\lambda)$.
Candidate hyperparameter sequences $\{\lambda^{(k)}_{i}:i\in\mathbb{N}^{+}\}$ for $k\in[K]$.
Weighting exponent $\xi$.
Output :
Online model selection index $\hat{k}^{*}_{i}$ at each time step $i$.
Initialize :
Set estimators $\hat{f}^{(1)}=\dots=\hat{f}^{(K)}=0$.
Set wRV statistics ${\rm RV}^{(1)}=\dots={\rm RV}^{(K)}=0$.
Set index $i=0$.
while there are remaining training samples do
       Read new sample $(X_{i+1},Y_{i+1})$for $k\in[K]$ do
             Predict $\hat{Y}_{i+1}^{(k)}\leftarrow\hat{f}^{(k)}(X_{i+1})$  Update wRV: ${\rm RV}^{(k)}\leftarrow{\rm RV}^{(k)}+i^{\xi}(Y_{i+1}-\hat{Y}_{i+1}^{(k)})^{2}$  Update estimator: $\hat{f}^{(k)}\leftarrow\text{Update}(\hat{f}^{(k)},(X_{i+1},Y_{i+1}),\lambda^{(k)}_{i+1})$
@@
      Output current selection: $\hat{k}_{i}^{*}\leftarrow\arg\min_{k}{\rm RV}^{(k)}$  Increment index: $i\leftarrow i+1$
@@
Algorithm 1 Stochastic approximation with weighted rolling validation (wRV) for regression.
The wRV statistics (9) is not only a natural target to examine when comparing online estimators, but also offers much computational savings when implemented efficiently. In Algorithm 1, we summarize our recommended framework on integrating model training and model selection. In contrast to the batch-learning scenario where training, validation, and testing are typically done in separated phases, we proposed a procedure that: 1) updates multiple estimators simultaneously; 2) calculates the wRV statistics interchangeably with model training; and 3) can offer a selected model at any time during the training process.
We take the kernel-SGD estimator (3) as an example to quantify the computational gain. Examining the update rule, we can verify that both the trajectory functions $\{\tilde{f}_{i}\}$ and the estimators $\{\hat{f}_{i}\}$ are a linear combination of $i$ kernel functions “centered” at the past sample covariate vectors. Formally, there are some $\tilde{\beta}_{mi},\hat{\beta}_{mi}\in\mathbb{R}:$
@@
Suppose evaluating one kernel function $\mathcal{K}(x,z)$ takes $O(p)$ computation—recall that $p$ is the dimension of $X_{i}$. Given a new sample $(X_{i+1},Y_{i+1})$, the majority of computation for the $\text{Update}(\cdot)$ operator is calculating $\mathcal{K}(X_{i+1},X_{m})$ for all $m\in[i]$. This means $O((p+C)i)$ calculation when processing one sample $(X_{i+1},Y_{i+1})$, where $C$ is a small constant corresponding to the extra computation after evaluating all $\mathcal{K}(X_{i+1},X_{m})$. Moreover, the above computational expense accumulates to the order of $O((p+C)n^{2})$ for processing the first $n$ samples $\{(X_{1},Y_{1}),...,(X_{n},Y_{n})\}$. Naively training $K$ kernel-SGD estimators one at a time with $n$ samples would require $O(K(p+C)n^{2})$ computation if no intermediate quantities are saved. However, when the $K$ sequences of trajectories only differ at the learning rate $\gamma_{i}$, they can be simultaneously updated with only $O((p+KC)n^{2})$ computation using the scheme in Algorithm 1, simply because the evaluated $\mathcal{K}(X_{i+1},X_{m})$ kernel value at each step can be shared between the estimators.
In addition to the computational gain for simultaneous training, these evaluated kernel values can be further shared with the wRV statistic calculation that only increases the total expense to at most $O((p+2KC)n^{2})$. This effectively makes performance assessment a “free lunch” relative to the cost of training, without a separate hold-out set for model selection. The only extra space expense is saving extra $K$ numbers $\{{\rm RV}^{(k)}\}$. A similar calculation also holds for sieve-SGD (4): the computational expense of training and validating $K$ estimators can be bounded by $O((p+2KC)nJ_{n})$, where $J_{n}$ is the largest basis number among the $K$ candidate estimators. For a fixed $p$, $J_{n}$ is typically a diverging sequence with $J_{n}=o(\sqrt{n})$.
@@ 4 Consistency of Rolling Validation, Fixed Number of Candidates
In this section, we analyze the statistical properties of the proposed RV procedure in regression settings. We focus on whether wRV can select $k^{*}$ specified in (8) from a finite candidate; the increasing candidate-cardinality case is covered in Section 5. We will show that under sufficient risk separation and stability of the wRV statistic, the proposed procedure can consistently select the superior model. For simplicity, we consider $K=2$ with two sequences of online estimators of $f_{0}$. This consistency result extends to any finite $K\geq 2$ by considering $k^{*}$ against each candidate and using the union bound. Below, we list the technical conditions.
@@
The data points $\{(X_{i},Y_{i})\in\mathbb{R}^{p+1},i\in\mathbb{N}^{+}\}$ are IID samples from a common distribution $P(X,Y)$. The centered noise variable $\epsilon_{1}=Y_{1}-f_{0}(X_{1})$ has finite variance: $\mathbb{E}[\epsilon_{1}^{2}|X_{1}=~{}x]\leq\sigma^{2}$ for any $x$ in the support of $X_{1}$.
@@
Next, we formalize the assumption on the sub-optimal model’s estimation accuracy, essentially stating that its average risk is no better than $Mi^{-a}$.
@@
$\liminf_{i\rightarrow\infty}i^{a}\mathbb{E}[r(\hat{f}_{i})]=M.$ (11)
Recall $r(\hat{f}_{i})=$ $\mathbb{E}_{X_{0}}\left[\left\{\hat{f}_{i}(X_{0})-f_{0}(X_{0})\right\}^{2}\right]$ and the outer expectation is unconditional, making the left-hand-side in (11) non-random. We impose a similar condition for the superior sequence, flipping the limit inferior to limit superior:
@@
$\limsup_{i\rightarrow\infty}i^{a}\mathbb{E}[r(\hat{f}_{i})]=M.$ (12)
@@
We assume $\sup_{i}\mathbb{E}[r(\hat{f}_{i})]<\infty$ and there exist constants $a\in[0,1]$, $0<C_{1}<\infty$ such that
$\limsup_{i\rightarrow\infty}i^{2a}\mathbb{E}\left[\left\{\hat{f}_{i}\left(X\right)-f_{0}\left(X\right)\right\}^{4}\right]=C_{1}.$ (13)
@@
Although 4.4 implies 4.3 with $M=C_{1}^{1/2}$, we state these conditions separately because, in Theorem 4.1, the superior and inferior sequences are assumed to satisfy different combinations of them. We will also compare the constants in (11) and (12), while the constant $C_{1}$ in (13) is of lesser importance. Additionally, we do not require the sequence $i^{a}\mathbb{E}[r(\hat{f}_{i})]$ itself to have a limit as $i\rightarrow\infty$.
@@
The values of $(M,a)$ reflect the quality of the estimators, which partially reflects the problem’s difficulty. They are determined by the fitting methods and tuning parameters. For optimal parametric SGD, $a=1$. For a minimax-optimal estimation procedure in a Sobolev ellipsoid (5)—as well as some related RKHS space—$a=2s/(2s+1)$. The constant $M$ is less explicit but usually depends on the axis lengths of the ellipsoid, interpreted as the Sobolev/RKHS norm of $f_{0}$ or as a measure of $f_{0}$’s smoothness. Mean-squared error (MSE) convergence results for online (nonparametric) estimators are available in the literature, including [10, 43, 40, 41, 9]. Specifically, [43] also provides an interesting lower bound for the MSE error.
In addition to the moment conditions above, we also need the following stability conditions to control certain “non-standard” deviation quantities that cannot be fully reduced to martingale-type concentration. Let $(X^{\prime}_{j},Y^{\prime}_{j})$ be an IID copy of $(X_{j},Y_{j})$ for a $j\geq 1$. The actual estimator $\hat{f}_{i}(x)$ is trained with samples $\bm{Z}_{i}=\{(X_{l},Y_{l}),l\in[i]\}$ and can be explicitly expressed as $\hat{f}_{i}(x)=\hat{f}_{i}(x;\bm{Z}_{i})$. We denote an imaginary dataset $\bm{Z}_{i}^{j}$, obtained by replacing $(X_{j},Y_{j})$ in $\bm{Z}_{i}$ by its IID copy $(X^{\prime}_{j},Y^{\prime}_{j})$. The imaginary estimator trained with $\bm{Z}_{i}^{j}$ is $\hat{f}_{i}(x;\bm{Z}_{i}^{j})$. We need conditions on the difference between $\hat{f}_{i}(x;\bm{Z}_{i})$ and its perturb-one version: $\nabla_{j}\hat{f}_{i}\left(x\right):=\hat{f}_{i}\left(x;\bm{Z}_{i}\right)-\hat{f}_{i}\left(x;\bm{Z}_{i}^{j}\right)$.
@@
$i^{2b}\mathbb{E}[\{\nabla_{j}\hat{f}_{i}(X)\}^{2}|F^{j}]\leq C_{2}\qquad\text{almost surely},$ (14)
where $F^{j}$ is the $\sigma$-algebra generated by the first $j$ samples in $\bm{Z}_{i}$.
@@
The sup-norm stability condition in (14) can be simplified and relaxed to $\mathbb{E}[\{\nabla_{j}\hat{f}_{i}(X)\}^{2}]=O(i^{-2b})$ if we strengthen the convergence condition in 4.4 to the sup norm: $\sup_{x}\{|\hat{f}_{i}(x)-f_{0}(x)|\}=O(i^{-a})$. In fact, we require the product $\mathbb{E}[\nabla_{j}\hat{f}_{i}(X)\{\hat{f}_{i}(X)-f_{0}(X)\}\mid F^{i}]$ to be small enough in the proof (equation (37)), which allows for Hölder’s-inequality type trade-off between the stability term and the estimation error.
@@
$\lim_{n\rightarrow\infty}\mathbb{P}\left[{\rm RV}\left(\left\{g_{i},i\in[n]\right\},\xi\right)<{\rm RV}\left(\left\{h_{i},i\in[n]\right\},\xi\right)\right]=1$